OpenStudy (anonymous):

which ordered pair is in the solution set of the system of equations, y = -x + 1 and y = x^2 + 5x + 6? A. (-5,-1) B. (-5,6) C. (5,-4) D. (5,2)

OpenStudy (anonymous):

you have a choice you can a) solve it or b) plug in each pair of numbers and see which one fits both equations what do you choose?

OpenStudy (anonymous):

hint: solving is probably easier

OpenStudy (ribhu):

also u can plot a graph

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

since \(y=y\) you can solve \[-x+1=x^2+5x+6\]

OpenStudy (anonymous):

add \(x\) subtract \(1\) and solve \[x^2+6x+5=0\] by factoring

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

6x^2 + 5 = 0 is what i got from factoring

OpenStudy (anonymous):

that is not factoring as a matter of fact, i am not sure what that is but \[x^2+6x+5\neq 6x^2+5\] in any way

OpenStudy (anonymous):

factoring means to write as a product, like \[x^2+2x+1=(x+1)(x+1)\] for example

OpenStudy (anonymous):

ah i see..

OpenStudy (anonymous):

can you factor \[x^2+6x+5\]? you do not have too many choices

OpenStudy (anonymous):

ok i am going to try one sec..

OpenStudy (anonymous):

(x+5)(6x+0) is the correct answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(6x+0=6x\) that is not right

OpenStudy (anonymous):

the \(x+5\) part is right

OpenStudy (anonymous):

i think the correct answer to the original question is D. (5,2)

OpenStudy (anonymous):

it is not

OpenStudy (anonymous):

damn, i should of paid attention to my teacher :(

OpenStudy (anonymous):

you really have to factor \(x^2+6x+5\) there is no avoiding it

OpenStudy (anonymous):

do you mind teaching me how to factor, like the steps?

OpenStudy (anonymous):

what two numbers multiply to get 5 and add to 6?

OpenStudy (anonymous):

3 and 2

OpenStudy (anonymous):

do you mean two numbers multiply to get 6, and add to get 5?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

5 and 1

OpenStudy (anonymous):

you have \[x^2+\color{red}6x+\color{blue}5\] two numbers that multiply to \(\color{blue}5\) and add to \(\color{red}6\)

OpenStudy (anonymous):

5 and 1 are two numbers that when multiplied make 5 and added make 6

OpenStudy (anonymous):

yeah 5 and 1 that means \[x^2+6x+5=(x+1)(x+5)\]

OpenStudy (anonymous):

if you multiply it out, you will see it is right now set each factor equal to zero and solve \[x+1=0\\ x=?\]

OpenStudy (anonymous):

you want me to solve the question above?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, what do you mean set each factor equal to zero? i dont understand what that means sorry

OpenStudy (anonymous):

the factors of \((x+1)(x+5)\) are \(x+1\) and \(x+5\) setting them equal to zero means... set them equal to zero \[\huge x+1=0\] solve for \(x\)

OpenStudy (anonymous):

oh i see, x + 1 = 0 and x + 5 = 0

OpenStudy (anonymous):

right now solve for \(x\)

OpenStudy (anonymous):

x + 1 = 0 x=1

OpenStudy (anonymous):

no \(1+1=2\) not \(1+1=0\) try again takes one step but if \(x+1=0\) then \(x\neq 1\)

OpenStudy (anonymous):

oh its x = -1

OpenStudy (anonymous):

i had a feeling it was -1, i had a feeling it was not right.

OpenStudy (anonymous):

the correct answer for the original question is A. (-5,-1) i see now! sweet!

OpenStudy (anonymous):

yes, that is it

OpenStudy (anonymous):

you're better than my math teacher .. TY!

OpenStudy (misty1212):

yw